Three friends buy a car. One pays 1/4 one pays 1/3 how much does the third pay?
Answer by nerdybill(7384)

Let x = fraction third person pays
then, since the sum has to total 1:
1/4 + 1/3 + x = 1
multiplying both sides by 12:
3 + 4 + 12x = 12
7 + 12x = 12
12x = 5
x = 5/12